A man invested $1000 by buying 100 shares of Go Big or Go Home Mutual Fund,...

  1. A man invested $1000 by buying 100 shares of Go Big or Go Home Mutual Fund, an aggressive growth no-load mutual fund. So far, the NAV for his investment has risen from $10 per share to $13.

What is the percentage increase in the NAV for this mutual fund?

If he redeemed the first 100 shares of his mutual fund investment for $13 per share, what would be his capital gain over the amount invested?
Answer #1

Percentage increase in NAV = (13 - 10)/10 = 30%

Capital Gain = (13 - 10)(100)

Capital Gain = $300